OpenStudy (anonymous):

this year 75% of the graduating class had taken at least 8 math courses. of the remaining class members, 60% had taken 6 or 7 math courses.What % of the graduating class had taken fewer than 6 math courses?

OpenStudy (anonymous):

suppose the total members of the class is x. 75% of x had taken at least 8 math courses. that means the remaining had taken fewer than 8 courses. the remaining class members is 100%x - 75%x = 25%x then 60% of the remaning class members had taken 6 or 7 math courses . since the remaining class members is 25% , teh class member that had taken 6 or 7 math courses is: 60% of 25%x = (60/100)(25x/100) = 15x/100 = 15%x so the class members that had taken fewer than 6 math courses is 25%x - 15%x = 10%x or 10% of the graduating class

OpenStudy (gw2011):

10% have taken fewer than 6 math courses. It is calculated as follows: 60% of the remaining 25% = 15%, and 75% + 15% = 90% which leaves 10% of the students taking fewer than 6 math courses.